# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v3

@@ 4.2 Above the Riemann height
Suppose $t_{0}\geq H$ and let there be a positive real number $d$. As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
$\sigma_{0}=1+\delta_{T_{0}}=1+\frac{d}{\log T_{0}},$ (34)
and extend their radii slightly into the critical strip, and then apply Theorem 1. This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros. We will instead utilise the zero-free region mentioned in Section 2.1 and keep our discs within said region. This requires discs with radii decreasing to 00 at the rate of $1/\log T_{0}$, similar to the zero-free region.
Let
$R_{T_{0}}=2\delta_{T_{0}}=\frac{2d}{\log T_{0}}\qquad\text{and}\qquad r=\alpha R_{T_{0}},\quad 0<\alpha<1.$ (35)
In addition, let $d$, and hence $R_{T_{0}}$, be chosen such that
$\sigma_{0}-R_{T_{0}}=\sigma_{0}-2\delta_{T_{0}}\geq 1-\frac{1}{W_{0}\log(T_{0}+R_{T_{0}})},$
which is fulfilled by
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R_{H}}{H}\big{)}}{\log H}\right)^{-1}.$ (36)
This ensures that the disc $|s-s_{0}|\leq R_{T_{0}}$ lies entirely in the zero-free region, so $f(s)$ is analytic there.
Next, we want to obtain a bound on $M$. This is done by choosing
$R^{\prime}=\frac{1}{2}-\delta_{H}=\frac{1}{2}-\frac{d}{\log H}$
in Lemma 2. Then as long as $d$ is chosen to satisfy (3), i.e.,
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R^{\prime}}{H}\big{)}}{\log H}\right)^{-1},$ (37)
we have for $\sigma\geq 1-\delta_{H}$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||.$ (38)
Estimating $M$ thus requires a uniform bound of $\zeta(s)$ in the region
$\frac{1}{2}\leq 1-\delta_{T_{0}}-R^{\prime}\leq\sigma\leq 1+3\delta_{T_{0}}+R^{\prime}\leq 1+\frac{1}{2}+2\delta_{H}.$
By Lemma 7 with $\delta_{r}=1/2+2\delta_{H}$, we have for $T_{0}\geq H$, and the region $1/2\leq\sigma\leq 3/2+2\delta_{H}$ that
$|\zeta(s)|\leq A_{2}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{3}T_{0}^{1/6}\log T_{0},$ (39)
where
$A_{3}$ $=A_{2}\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right),$ (40)
with
$A_{2}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6}$ (41)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right).$
Combining (39), (41), and (40) with (38) thus gives for $1-\delta_{T_{0}}\leq\sigma\leq 1+3\delta_{T_{0}}$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{3}\right).$ (42)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we use the trivial bound for $\sigma>1$, given in (31). This yields
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq\frac{1}{\sigma_{0}-1}\leq\frac{\log T_{0}}{d}.$ (43)
Up to this point, Theorem 1 will give a result that holds for
$1-\delta_{T_{0}}(2\alpha-1)\leq\sigma\leq 1+\delta_{T_{0}}(2\alpha+1),$ (44)
and for a result on any larger $\sigma$, we simply use (31). At the same time, if we want a result that holds for $\sigma\geq 1-1/W\log T_{0}$, then by the left inequality of (44), we must have
$W=\frac{1}{d(2\alpha-1)}\qquad\text{and}\qquad\alpha=\frac{1}{2}+\frac{1}{2dW}.$ (45)
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
@@
Let $d$ and $W>W_{0}$ be positive parameters, subject to (36), (37), and (45). Let also $10<\alpha<$ 1 be a parameter defined in (45), and depending on $d$ and $W$. In a region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H$
@@
$Q_{H}=\max\left\{\frac{\lambda}{6}+\frac{1+\alpha}{d(1-\alpha)}+\frac{\lambda(\log\log H+\log A_{3})}{\log H},\,\frac{1}{d(1+2\alpha)}\right\},$
with $A_{3}$ defined in (40),
$\lambda=\frac{8\alpha}{(1-\alpha)\pi R^{\prime}},\qquad\text{and}\qquad R^{\prime}=\frac{1}{2}-\frac{d}{\log H}.$
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
For the case of $t\geq H$, with a fixed choice of $W$ as defined in (45), it remains to optimise $Q_{H}$ in Theorem 3 over $\alpha$ and $d$, subject to the conditions mentioned in said theorem, while also checking that (11) holds. We found that the optimal value for $d$ in each case was approximately $d=0.179898\ldots$, and any differences in the final result caused by varying $d$ further occurred only after the sixth decimal place. We note also that the optimal $d$ is one that is as close to $1/W_{0}$ as possible while still satisfying (36) and (37).
On the other hand, a simple calculation shows that
$0.821-\frac{1}{W_{0}\log t}>$ 0.82
for $t\geq t_{0}=e$. Now, for $\sigma\geq 0.8$ and $5\leq t\leq H$, Corollary 3 and Table 2 asserts that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=33.249\log t,$
@@
First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $5\leq t\leq H$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=9.161\log t.$
Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha$ even further, since we now require $\sigma_{0}-\alpha R_{T_{0}}\geq 1$, which is equivalent to $\alpha\leq 1/2$. In order for the result to hold for all $\sigma\geq 1$, we choose $\alpha=1/2$, so in this case,
$Q_{H}=\max\left\{\frac{8}{6\pi R^{\prime}}+\frac{3}{d}+\frac{8(\log\log H+\log A_{3})}{\pi R^{\prime}\log H},\,\frac{1}{2d}\right\}.$
As before, we then optimise over $d$, while ensuring that we fulfil the conditions in Theorem 3 and (11). The optimal value found was $d=0.179898$. ∎
$W$ $Q$ $\alpha$ $W$ $Q$ $\alpha$
$5.559$ $450041$ $0.999972$ $5.59$ $4458$ $0.997200$
$5.56$ $106223$ $0.999882$ $5.6$ $3383$ $0.996312$
$5.5601$ $98674$ $0.999873$ $5.7$ $1002$ $0.987604$
$5.5602$ $92137$ $0.999864$ $5.8$ $594.087$ $0.979198$
$5.5603$ $86414$ $0.999855$ $5.9$ $425.318$ $0.971076$
$5.5604$ $81354$ $0.999846$ $6$ $333.034$ $0.963224$
$5.5605$ $76859$ $0.999837$ $6.1$ $274.847$ $0.955630$
$5.561$ $60223$ $0.999792$ $6.2$ $234.806$ $0.948282$
$5.562$ $42026$ $0.999703$ $6.3$ $205.568$ $0.941166$
$5.563$ $32278$ $0.999613$ $6.4$ $183.281$ $0.934273$
$5.564$ $26201$ $0.999523$ $6.5$ $165.729$ $0.927592$
$5.565$ $22051$ $0.999433$ $6.75$ $134.741$ $0.911755$
$5.566$ $19036$ $0.999343$ $7$ $114.502$ $0.897049$
$5.567$ $16748$ $0.999254$ $8$ $74.999$ $0.847418$
$5.568$ $14951$ $0.999164$ $9$ $58.454$ $0.808816$
$5.569$ $13502$ $0.999074$ $10$ $49.359$ $0.777935$
$5.57$ $12310$ $0.998985$ $11$ $43.607$ $0.752668$
$5.575$ $8541$ $0.998537$ $12$ $39.642$ $0.731612$
$5.58$ $6542$ $0.998091$ $13$ $36.742$ $0.713796$
Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and $\alpha$ (rounded to six decimal places). Each entry is valid for $t\geq 5$.
@@
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [19]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [7, Lemma 10]).
However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. Furthermore, through a trick utilising the trigonometric polynomial, we are able to get an slight asymptotic improvement on the bounds, at the cost of a slightly worse constant. As before, the proofs will be split into two cases, $t\leq H$ and $t\geq H$.
5.1 Assuming a partial Riemann hypothesis
@@
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{RH(T)}(d,d_{1},\sigma_{1},t_{0},\eta)\log t,$
@@
$Y_{RH(T)}$ $(d,d_{1},\sigma_{1},t_{0},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\Bigg{\}},$
with
$A_{4}$ $=\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)^{7/6},$
$B_{t_{0}}$ defined in Theorem 2, and $V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta),\,\eta$ defined in Lemma 12.
@@
Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Since by $RH(T)$ there are no zeros for $\sigma\geq 1/2$, we do not have to check the condition in (3). The same argument in Section 4.1 with Lemma 2 leading to (30), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},$
and in return gives us
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Thus we have from (46) that

### v4

@@ 4.2 Above the Riemann height
$W$ $Q$ $\beta$ $\epsilon_{1}$ $\sigma_{1}$ $\eta$
$5.559$ $928465$ $0.999980$ $0.012554$ $1.756408$ $1.050104$
$5.56$ $170199$ $0.999890$ $0.012551$ $1.937703$ $2.228745$
$5.5601$ $157356$ $0.999881$ $0.012563$ $1.393177$ $4.286069$
$5.5602$ $146346$ $0.999872$ $0.012587$ $1.710791$ $1.367807$
$5.5603$ $136750$ $0.999863$ $0.012582$ $1.776753$ $1.141731$
$5.5604$ $128379$ $0.999854$ $0.012712$ $1.889117$ $2.700506$
$5.5605$ $120884$ $0.999845$ $0.012614$ $1.711461$ $3.302948$
$5.561$ $93736$ $0.999800$ $0.012690$ $1.339371$ $2.627981$
$5.562$ $64686$ $0.999710$ $0.012842$ $1.299159$ $2.599102$
$5.563$ $49364$ $0.999620$ $0.012699$ $1.427545$ $1.952944$
$5.564$ $39919$ $0.999531$ $0.012747$ $1.557141$ $2.691489$
$5.565$ $33530$ $0.999441$ $0.013347$ $1.715049$ $4.337899$
$5.566$ $28887$ $0.999351$ $0.013295$ $1.234376$ $4.393614$
$5.567$ $25383$ $0.999262$ $0.013634$ $1.973530$ $3.744506$
$5.568$ $22625$ $0.999172$ $0.013225$ $1.814250$ $2.845626$
$5.569$ $20414$ $0.999082$ $0.013180$ $1.641377$ $1.985693$
$5.57$ $18592$ $0.998993$ $0.012868$ $1.322806$ $4.407476$
$5.575$ $12870$ $0.998545$ $0.012963$ $1.920818$ $3.555767$
$5.58$ $9863$ $0.998102$ $0.019214$ $1.690970$ $2.264745$
$5.59$ $6698$ $0.997207$ $0.012557$ $1.929368$ $4.355523$
$5.6$ $5082$ $0.996321$ $0.015401$ $1.421225$ $3.490106$
$5.7$ $1501$ $0.987620$ $0.025973$ $1.590960$ $3.434482$
$5.8$ $888.269$ $0.979209$ $0.018833$ $1.305491$ $4.203009$
$5.9$ $635.099$ $0.971088$ $0.021934$ $1.977612$ $4.803605$
66 $496.670$ $0.963240$ $0.027122$ $1.420679$ $3.567866$
$6.1$ $409.353$ $0.955640$ $0.016196$ $1.713398$ $1.710689$
$6.2$ $349.287$ $0.948293$ $0.019949$ $1.623774$ $3.117029$
$6.3$ $305.581$ $0.941204$ $0.066228$ $1.701866$ $3.427829$
$6.4$ $272.026$ $0.934286$ $0.023358$ $1.163604$ $3.125341$
$6.5$ $246.170$ $0.927735$ $0.255900$ $1.454736$ $3.093587$
$6.75$ $199.284$ $0.911787$ $0.057471$ $1.361313$ $4.269028$
77 $168.924$ $0.897081$ $0.059017$ $1.777110$ $3.321819$
88 $109.668$ $0.847438$ $0.038048$ $1.707618$ $2.623961$
99 $84.858$ $0.808831$ $0.031281$ $1.249275$ $2.976333$
$10$ $71.220$ $0.777942$ $0.016334$ $1.624690$ $4.127955$
$11$ $62.611$ $0.752706$ $0.084730$ $1.177847$ $1.792189$
$12$ $56.653$ $0.731621$ $0.020106$ $1.606738$ $2.688990$
$13$ $52.306$ $0.713814$ $0.041793$ $1.671118$ $3.367414$
Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and parameters $\beta$, $\epsilon_{1}$, $\sigma_{1}$, $\eta$ (all rounded to six decimal places). Each entry is valid for $t\geq 13$.
@@
Let $d\leq 1/W_{0}$, $\epsilon_{1}\leq 1/2$, and $\beta<1$ be positive real parameters, subject to (35) and (38). Define $W>W_{0}$ by
$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$ (30)
where the function $a_{\epsilon_{1}}(t_{0})$ also depends on $\epsilon_{1}$ and is defined in (34).
Then in a region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq t_{0}\geq H$
@@
$Q_{H}=\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d)}\right\},$
with
$\lambda_{1}=\frac{16\beta}{(1-\beta)}\left(1-\frac{8d}{2d+\log t_{0}}\right)^{-2}\qquad\text{and}\qquad\lambda_{2}=\frac{1+\beta}{d(1-\beta)},$ (31)
provided that
$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\frac{2d}{\log t_{0}})}{0.618(1+a_{1}(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}))^{1/6}}$ (32)
is satisfied.
The quantity $A_{3}$ is defined in (4.2), and depends on $d$ and $t_{0}$, on $a_{0}$ and $a_{1}$, which are defined as in Lemma 4, and on $V(\kappa,\sigma_{1},t_{0},\eta)$ and its corresponding parameters, defined as in Lemma 12 and subject to (37).
Proof.
The proof is similar to that of Theorem 2. Suppose $t_{0}\geq H$ and let there be a positive real parameter $d\leq 1/W_{0}$ to be determined later. As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
$\sigma_{0}=1+\frac{d}{\log T_{0}},\qquad r=\frac{1}{2}+\frac{d}{\log T_{0}},$ (33)
extend their radii slightly into the critical strip, and then apply Lemma 2 with $f(s)=\zeta(s)$, $s=\sigma+iT_{0}$, and $A_{1}$ defined by (25). This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros. We will instead utilise the zero-free region mentioned in §2.1 and keep our discs within said region. This requires discs with radii decreasing to 0 at the rate of $1/\log T_{0}$, similar to the zero-free region.
Let $\epsilon_{1}$ be a fixed positive real parameter, and for the sake of brevity define the decreasing function (for $T_{0}>1$)
$a_{\epsilon_{1}}(T_{0}):=1+\frac{\log\left(1+\frac{\epsilon_{1}}{\log T_{0}}\right)}{\log T_{0}}>1.$ (34)
To ensure that the non-vanishing condition on $f(s)=\zeta(s)$ is met, notice that if we impose the condition $\alpha r\leq\epsilon_{1}\leq r$, then the disc $|s-s_{0}|\leq\alpha r$ with
$\alpha r\leq\sigma_{0}-\left(1-\frac{d}{\log(T_{0}+\epsilon_{1})}\right),$
is completely contained in the zero-free region as defined in §2.1. This also implies that the following choice of $\alpha$ and bounds are admissible:
$\alpha:=\frac{2d\left(1+a_{\epsilon_{1}}(t_{0})^{-1}\right)}{\log T_{0}+2d}<\frac{4d}{2d+\log t_{0}}\qquad\text{and}\qquad\frac{2d}{\log t_{0}}\leq\epsilon_{1}\leq\frac{1}{2},$ (35)
which in turn requires that $d<(\log t_{0})/6$ since we need $\alpha<1/2$ for Lemma 2. This is easily achieved since $d\leq 1/W_{0}<(\log t_{0})/6$ for all $t_{0}\geq 3$.
Next, we determine $A_{1}$ and $A_{2}$. Suppose (6) is satisfied with $1+\delta_{r}=\sigma_{0}+r$ and define $A_{1}$ (which depends on $A_{3}$) by (25), but use (33) with the argument of Lemma 12 to replace (4.1) by
$A_{3}:=0.618\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)^{7/6}$
$\bm{\cdot}\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)}{\log t_{0}}\right)V\left(\frac{d}{\log t_{0}},\sigma_{1},t_{0},\eta\right),$ (36)
where $\sigma_{1}$ and $\eta$ satisfies
$1+\frac{d}{\log t_{0}}\leq\sigma_{1}\leq 2\qquad\text{and}\qquad 0<\eta<2t_{0}.$ (37)
On the other hand, $A_{2}$ is determined by invoking (28), which gives us
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq A_{2}:=\frac{1}{d}\log T_{0}.$
Now, for any results from Lemma 2 to be meaningful for us, we require the left most point of $|s-s_{0}|\leq\alpha\beta r$ to lie within the zero-free region and in the critical strip. In other words, $\sigma_{0}-\alpha\beta r\leq 1$, which implies
$\beta\geq\frac{2d}{\alpha(2d+\log T_{0})}\geq\left(1+\frac{1}{a_{\epsilon_{1}}(t_{0})}\right)^{-1}\geq\frac{1}{2}.$ (38)
In particular, since
$\sigma_{0}-\alpha\beta r=1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}},$
to obtain a result that holds for
$1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}}\leq\sigma\leq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)\frac{1}{\log T_{0}},$
we choose $W>W_{0}$ such that
$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$
while to cover the remaining range of $\sigma\geq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)/\log T_{0}$, we again use (28), and take the maximum of the two results.
Finally, putting everything together into Lemma 2 and making the substitution $T_{0}\mapsto t$, we prove our result. ∎
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
For the case of $t\geq H$, with a fixed choice of $W$ as defined in (30), it remains to optimise $Q_{H}$ in Theorem 3 over $d$, $\beta$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, subject to the conditions mentioned in said theorem, while also checking that (32) holds, which is done in a similar way as in the proof of Corollary 3. We found that the optimal value for $d$ in all cases is $d=1/W_{0}$, and (32) holds for all $t\geq t_{0}\geq H$. All computed parameters are presented in Table 3, rounded to six decimal places.
On the other hand, a simple calculation shows that for $t\geq t_{0}=13$,
$1-\frac{1}{W_{0}\log t}>0.929.$
Now, for $\sigma\geq 0.9$ and $13\leq t\leq H$, Corollary 3 and Table 2 asserts that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=13.123\log t,$
@@
First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $13\leq t\leq H$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=8.101\log t.$ (39)
Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha\beta r$ even further, since we can now set $\sigma_{0}-\alpha\beta r=1$, which is equivalent to
$\beta=(1+a_{\epsilon_{1}}(t_{0})^{-1})^{-1}.$ (40)
So in this case, for $\sigma\geq 1$ and $t\geq t_{0}\geq H$, we have that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{2d}\right\}\,\log t,$ (41)
where $\lambda_{1}$ and $\lambda_{2}$ are defined as in (31), but with the added restriction of (40).
Similar to the proof of Corollary 1, we then optimise (41) over $d$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, while ensuring that we fulfil the appropriate conditions in Theorem 3 and (32). The optimal value of $d$ found was again $d=1/W_{0}$. We then compare the result with (39) and take the maximum to obtain a bound that holds for all $t\geq 13$. ∎
@@
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [20]), while the above integral is estimated by bounding the real part of the logarithmic derivative by a uniform bound on its absolute value $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by using the method of [7, Lemma 10]), which splits the interval $\sigma,1+\delta_{1}]$ and uses the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval.
Furthermore, through an argument utilising the trigonometric polynomial, we are able to get a slight asymptotic improvement on the bounds, at the cost of a worse constant.
@@
Let $t_{0}\geq 13$ and $d_{1}>0$ be a constant. Let $\{W_{j}:1\leq j\leq J\}$ be a sequence of increasing real numbers where $(W_{j},Q_{1,j})$ is a pair for which $t\geq t_{0}$ and
$\sigma\geq 1-\frac{1}{W_{j}\log t}\quad\text{implies}\quad\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{1,j}\log t.$
Then for
$\sigma\geq 1-\frac{1}{W_{1}\log t},$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{0}(d_{1},\sigma_{1},t_{0},\eta)\log t,$
@@
$Y_{0}$ $(d_{1},\sigma_{1},t_{0},\eta)=\max\Biggl{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$ (43)
$V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\cdot\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\Biggr{\}},$
with any $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$, $\gamma$ denoting Euler’s constant, and $V$ defined as in Lemma 12.
@@
Let $\delta_{1}=\delta_{1}(t):=d_{1}/\log t$. We will split the proof into two cases: $\sigma\leq 1+\delta_{1}$ and $\sigma>1+\delta_{1}$. First see that in the range $1-1/W_{1}\log t\leq\sigma\leq 1+\delta_{1}$ we have